# End-to-End 4D Heart Mesh Recovery Across Full-Stack and Sparse Cardiac MRI

### v1

@@ 3.3 Initial Heart Model Reconstruction
Given the $F^{0}_{\rm static}$ features computed as discussed above from a full-slice CMR image $\mathbf{I}^{0}$, our first task is to reconstruct the initial tetrahedral mesh $\mathcal{G}^{0}$. To this end, we start from an initial tetrahedral grid $\mathcal{G}^{init}$ obtained by uniformly sampling a unit cube. We then trilinearly interpolate $F^{0}_{\rm static}$ at the vertices of $\mathcal{G}^{init}$. An SDF value and an offset for each vertex is then predicted using a GCN to create $\mathcal{G}^{0}$. We write
@@
where $\mathbf{v}_{i}$ is a grid vertex of $\mathcal{G}^{init}$ and $\mathbf{v}_{i}^{\prime}$ the same grid vertex after deformation. $F^{0}_{\rm static}(\mathbf{v})$ denotes feature extraction at $\mathbf{v}$ from $F^{0}_{\rm static}$ with trilinear interpolation, and $\cdot$$\cdot$] denotes concatenation. To enhance the network’s spatial sensitivity, the vertex location is appended to the feature vector before it is passed to the GCN. For datasets with $C$ classes, we simply set the output channel of the GCN layer to $C\times 4$, 3 for deformation, and 1 for SDF value. This yields a separate tetrahedral mesh for each class. Following prior works [43, 44, 45, 46], our network also has an auxiliary segmentation prediction branch during training time.
@@ 3.4 Motion Recovery
In Eq. 5, the vertex features are extracted by trilinear interpolating $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$ at the vertex locations. They are then concatenated and passed through Eq. 6-8 to gradually update the vertex position, which aggregates the spatial information. Eq. 5-8 is repeated $K$ times, which we set to $2$ by default.
@@ 3.5 Two-Stage Weakly Supervised Training Using Only Keyframe Annotations
Creating high-quality annotations for an entire sequence is costly. As a result, many datasets provide annotations only for keyframes, such as end-diastolic (ED) and end-systolic (ES) phases. This rules out full supervision. To address this, we propose a two-stage weakly-supervised training scheme which makes training the reconstruction and motion branches feasible using only the available keyframe annotations.
3.5.1 Training the Reconstruction Branch
We start by training the reconstruction branch of Section 3.3. At each training iteration, we randomly select a labeled image $\mathbf{I}$ with ground truth segmentation and mesh annotation $L^{gt},\mathcal{M}^{gt}$. Let $L^{p},\mathcal{G}^{p}$ be our model’s segmentation and tetrahedra prediction. We minimize the loss
@@ 3.5.1 Training the Reconstruction Branch
3.5.2 Training the Deformation Estimator
We then train the motion branch. At each iteration, we randomly select a sequence $\{\mathbf{I}^{t}\}_{t=0}^{T}$, and then choose an unlabeled frame $\mathbf{I}^{u}$ and a labeled frame $\mathbf{I}^{l}$ from it. We first predict $\mathcal{G}^{u}$ from $\mathbf{I}^{u}$. Next, assuming that $\mathbf{I}^{l}$ comprises a total of $D$ slices, we randomly pick 1 to $D$ 2D slices at random location to form $\hat{\mathbf{I}}^{l}$ to simulate the online few-slice case. $V_{\text{motion}}^{l}$ and $\hat{V}_{\text{motion}}^{l}$ are generated from $\mathbf{I}^{l}$ and $\hat{\mathbf{I}}^{l}$ using the AFA module. We write
$V_{\text{motion}}^{l}=AFA(F^{u}_{2d},F^{l}_{2d}),\hat{V}_{\text{motion}}^{l}=AFA(F^{u}_{2d},\hat{F}^{l}_{2d})\;.$ (10)
@@ 3.5.2 Training the Deformation Estimator
where $\mathcal{M}^{l}$ is the ground truth mesh for $\mathbf{I}^{l}$ and sg denotes the stop-gradient operation. Minimizing the distillation loss $\mathcal{L}_{distill}$ ensures that the 3D information from full-slice $\mathbf{I}^{l}$ is successfully distilled into the few-slice encoded feature, while minimizing the chamfer losses enables the deformation model to infer accurate motion from the feature map encoded from any number of 2D slices, while retaining the ability to infer motion from complete 3D full-slice feature map. As a result, minimizing $\mathcal{L}_{motion}$ guarantees good reconstruction results for both full and few-slice inputs.
Once trained, our model can be used for new patients without retraining. Note that our training strategy actually enables the model to predict cardiac deformation starting from any phase, not just the end-diastolic (ED) phase. Additionally, by employing an aggressive sampling scheme, the model learns to reconstruct motion from a highly diverse set of input configurations. As a result, even though the online scenario is simulated during training, our model demonstrates strong generalization capability. It can robustly handle cases where $\mathbf{O}^{t}$ originates from a different sequence or where there is a scanning angle discrepancy relative to $\mathbf{I}^{0}$, consistently producing reliable motion predictions despite these challenges as will be shown in the experiment section.
@@ 4 Experiments
4.1 Experimental Settings
Table 1: Few-slice quantitative evaluation on unified dataset. We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes. 1- / 5-Slice: Number of slices used to recover the motion. Due to space limit, we omit the standard deviation.
@@ 4.1 Experimental Settings
4.1.1 Data
We train and validate our method on a unified dataset constructed from three large and diverse publicly available datasets: ACDC, M&Ms, and M&Ms-2 [19, 20, 21]. ACDC features CMR sequences of about 30 frames for 100 patients, each one covering at least one cardiac cycle and with a slice thickness of 5-10 mm. For these datasets, segmentation masks for the left ventricle (LV), right ventricle (RV), and myocardium (MYO) are provided, but only for the end-diastolic (ED) and end-systolic (ES) frames. M&Ms and M&Ms-2 are made of multi-center and multi-vendor collected CMR sequences for 375 and 360 subjects, respectively. While featuring similar resolutions and sequence lengths as ACDC dataset, these two datasets cover a broader range of cardiac pathologies. As in earlier studies [13, 14, 9, 3], we register and fit a Statistic Shape Model (SSM) [24, 5] to segmentation mask at the ED and ES frames using non-rigid image registration methods to generate cardiac meshes. The unified dataset comprises 835 CMR sequences, randomly split into 501 (60%), 167 (20%), and 167 (20%) sequences for training, validation, and testing, respectively.
We use two additional datasets for external evaluation to assess generalization in a clinical context where model retraining is not feasible. The first dataset was collected at the Center for Interventional MRI, University Hospital Lausanne (CHUV), designed with the target online scenario in mind and hereafter referred to as CHUV. For each subject, a stack of short-axis (SAX) breath-hold cine slices was acquired at rest on a 1.5T interventional system (Siemens Healthineers, Magnetom Aera) using an accelerated steady-state free precession (SSFP) sequence covering the entire left ventricle. In addition, single SAX cine slices were acquired at mid-ventricular levels with slight angulations relative to the full SAX stack. Because both heart shape and rate may vary during intervention, updating of the 3D heart model is often required to ensure precise intracardiac manipulations. To mimic heart-rate changes, five volunteers exercised inside the interventional MRI (bending and extending their legs in two parallel plastic gutters for 2–3 minutes), thereby increasing their heart rates by approximately 20–30 beats per minute. Once elevated, single SAX cine slices were reacquired at the same mid-ventricular levels as during rest. For real-time acquisition, the online sequences had lower spatial resolution than the offline short-axis CMR stacks. To evaluate full-sequence performance across the entire cardiac cycle beyond the keyframes, we further use the second external dataset, 4DM [14], which focuses on left ventricular myocardium reconstruction and provides mesh annotations for every frame. This dataset enables a rigorous evaluation of the proposed weakly supervised motion learning scheme.
4.1.2 Evaluation Metrics
@@
For evaluation, we compare the predicted 3D mesh obtained by deforming the ED frame to the ES frame with the ground-truth 3D mesh at the ES frame. We use L2-norm Chamfer Distance (CD) to measure 3D reconstruction quality. Dice score is also reported to quantify the intra-slice segmentation results. Besides these quality measures, we also report the Mean-Absolute Error (MAE) of LV/RV ES Volume (LVESV/RVESV), LV/RV ejection fraction (LVEF/RVEF) derived from our predicted meshes. LVEF and RVEF are representative cardiac function indexes which could be used for clinical diagnosis.
4.1.3 Implementation Details
4.1.4 Baselines
We compare our method against several existing approaches for 4D shape reconstruction. For the online interventional setting, as noted earlier, most methods assume access to a complete CMR stack, making them unsuitable for the online few-slice setting. Therefore, we adapted several baselines so that they can operate under sparse inputs. 4DMR [14] is an implicit representation method that reconstructs myocardium mesh sequence by optimizing decoupled shape and motion latent codes at test time, minimizing the discrepancy between the predicted mesh and ground-truth contour points. We adapt it by restricting the optimization to incomplete contours extracted from the available few slices, while still using full contours for the initial frame. MR-Net [26] is a template-mesh-based approach that takes ground-truth contour points as input and leverages a hybrid PointNet + 3D CNN architecture to predict vertex deformations. Since MR-Net directly encodes point sets, it can naturally accommodate incomplete inputs, a scenario already considered in the original paper. Although originally designed for static reconstruction, we adapt it for motion estimation by setting the ED-frame mesh as the template.
@@ 4.1.4 Baselines
Figure 2: 1-Slice Comparison. (Left) We predict 3D meshes by deforming from ED to ES frame. Color indicates the magnitude of point-to-surface error in $m$. (Right) Segmentation results on slices at the apex and base location. Note that 4DMR can only output motion for myocardium.
@@
Figure 3: Motion sequence prediction results using different number of slices by deforming from ED frame to the target frame. Corresponding Volume-frame curve is given on the right. #1: NOR, #2: DCM.
@@ 4.2 Motion Estimation from a Few Slices
Table 4: Zero-shot evaluation results on public 4DM dataset. Check marks denote the dataset used for training. Oracle is trained and evaluated on 4DM.

### v2

@@ 3.3 Initial Heart Model Reconstruction
Given $F^{0}_{\rm static}$ computed as discussed above, our first task is to reconstruct the initial tetrahedral mesh $\mathcal{G}^{0}$. To this end, we start from tetrahedral grid $\mathcal{G}^{init}$ obtained by uniformly sampling a unit cube. We then trilinearly interpolate $F^{0}_{\rm static}$ at the vertices of $\mathcal{G}^{init}$. An SDF value and an offset for each vertex is predicted using a GCN to create $\mathcal{G}^{0}$. We write
@@
where $\mathbf{v}_{i}$ is a grid vertex of $\mathcal{G}^{init}$ and $\mathbf{v}_{i}^{\prime}$ the same grid vertex after deformation. $F^{0}_{\rm static}(\mathbf{v})$ denotes feature extraction at $\mathbf{v}$ from $F^{0}_{\rm static}$ with trilinear interpolation, and $\cdot$$\cdot$] denotes concatenation. To enhance spatial sensitivity, the vertex location is appended before it is passed to the GCN. For datasets with $C$ classes, we set the output channel to $C\times 4$, 3 for deformation and 1 for SDF. Following prior works [46, 21, 7, 12], our network also has an auxiliary segmentation prediction branch during training time.
@@ 3.4 Motion Recovery
In Eq. 3, the vertex features are extracted by trilinear interpolating $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$ at their locations. They are then concatenated and passed through Eq. 4-6 to gradually update the vertex position. Eqs. 3-6 are repeated twice.
@@ 3.5 Two-Stage Weakly Supervised Training Using Only Keyframe Annotations
Creating high-quality annotations for an entire sequence is costly. Therefore, many datasets provide annotations only for keyframes, such as end-diastolic (ED) and end-systolic (ES) phases. This rules out full supervision. To address this, we propose a two-stage weakly-supervised pipeline, which makes training the reconstruction and motion branches feasible using only the available keyframe annotations.
Training the Reconstruction Branch. We start by training the reconstruction branch of Section 3.3. At each training iteration, we randomly select a labeled image $\mathbf{I}$ with ground truth segmentation and mesh annotation $L^{gt},\mathcal{M}^{gt}$. Let $L^{p},\mathcal{G}^{p}$ be our model’s segmentation and tetrahedra prediction. We minimize the loss
@@
Training the Deformation Estimator. We then train the motion branch. At each iteration, we randomly select a sequence $\{\mathbf{I}^{t}\}_{t=0}^{T}$, and choose an unlabeled frame $\mathbf{I}^{u}$ and a labeled frame $\mathbf{I}^{l}$ from it. We first predict $\mathcal{G}^{u}$ from $\mathbf{I}^{u}$. Next, assuming that $\mathbf{I}^{l}$ comprises a total of $D$ slices, we randomly pick 1 to $D$ 2D slices at random location to form $\hat{\mathbf{I}}^{l}$ to simulate the online few-slice case. $V_{\text{motion}}^{l}$ and $\hat{V}_{\text{motion}}^{l}$ are generated from $\mathbf{I}^{l}$ and $\hat{\mathbf{I}}^{l}$ using the AFA module. We write
$V_{\text{motion}}^{l}=AFA(F^{u}_{2d},F^{l}_{2d}),\hat{V}_{\text{motion}}^{l}=AFA(F^{u}_{2d},\hat{F}^{l}_{2d})\;.$
@@
where $\mathcal{M}^{l}$ is the ground truth mesh for $\mathbf{I}^{l}$ and sg denotes the stop-gradient operation. Minimizing the distillation loss $\mathcal{L}_{distill}$ encourages features from few-slice input $\hat{\mathbf{I}}^{l}$ to align with the one from full-slice input $\mathbf{I}^{l}$, while minimizing the chamfer losses enables the deformation model to infer accurate motion from the feature map encoded from any number of 2D slices, while retaining the ability to infer motion from complete 3D full-slice feature map. As a result, minimizing $\mathcal{L}_{motion}$ guarantees good reconstruction results for both full and few-slice inputs.
Once trained, our model can be used for new patients without retraining. By employing an aggressive sampling scheme, the model learns to reconstruct motion from a highly diverse set of input configurations. As a result, even though the online scenario is simulated during training, our model demonstrates strong generalization capability. It can robustly handle cases where $\mathbf{O}^{t}$ originates from a different sequence or where there is a scanning angle discrepancy relative to $\mathbf{I}^{0}$, consistently producing reliable motion predictions as will be shown in the experiment section.
@@ 4 Experiments
Table 1: Few-slice quantitative evaluation on unified dataset. We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes. 1- / 5-Slice: Number of slices used to recover the motion.
@@
4.1 Experimental Settings
Datasets and Metrics. We train and validate our method on a unified dataset constructed from three large and diverse publicly available datasets: ACDC, M&Ms, and M&Ms-2 [5, 9, 25]. Segmentation for the left ventricle (LV), right ventricle (RV), and myocardium (MYO) are provided, but only for the ED and ES frames. The unified dataset comprises 835 CMR sequences, randomly split into training, validation, and testing sets in a ratio of 60%, 20%, and 20%. Dataset details are provided in Appendix A.1.
We use two additional datasets for external evaluation to assess generalization in a clinical context where model retraining is not feasible. The first dataset, referred to as iMRI, was collected on an interventional MRI system under both rest and elevated-heart-rate conditions. It is designed with the target online scenario in mind. For each subject, full short-axis cine stacks were acquired at rest, while additional single mid-ventricular slices were collected at rest and during mild in-scanner exercise to simulate heart-rate variations. The dataset thus reflects realistic cardiac motion dynamics encountered in online interventional MRI. Details on the scanning protocol, equipment, and exercise procedure are provided in Appendix A.2.
To evaluate full-sequence performance across the entire cardiac cycle beyond the keyframes, we use the second external dataset, 4DM [52], which focuses on left ventricular myocardium reconstruction and provides mesh annotations for every frame. This dataset allows rigorous evaluation of the proposed weakly supervised motion learning scheme.
@@ 4.1 Experimental Settings
For evaluation, we use the L2-norm Chamfer Distance (CD) to measure 3D reconstruction quality and the Dice score to quantify intra-slice segmentation results. We also report the Mean-Absolute Error of LV/RV ES Volume (LVESV/RVESV), LV/RV ejection fraction (LVEF/RVEF) derived from our predicted meshes. LVEF and RVEF are representative cardiac function indexes used for clinical diagnosis. See Appendix A.3 for details.
Implementation Details. In all experiments, the slices are resampled by linear interpolation to a spacing of $1.25\times 1.25$ mm. From this, the AFA module constructs 3D volume with an effective spacing of $1.25\times 1.25\times 2$ mm. See Appendix B for model and optimization details.
Baselines. We compare our method against several existing approaches for 4D shape reconstruction. For the online interventional setting, as noted earlier, most methods assume access to a complete CMR stack, making them unsuitable for few-slice inputs. Therefore, we adapted several baselines to operate under sparse inputs, including modified versions of 4DMR [52] and MR-Net [10] We also include two variants derived from shape reconstruction methods: MeshDeformNet [21] and DeepCSR [12]. We denote them as Ours-Mesh and Ours-SDF since their main differences with ours lie in the shape representation. For the traditional offline setting, in addition to the above methods, we further compare against Free Form Deformation (FFD[38], diffeomorphic Demons (dDemons[44], MR-Net w. nnU-Net, MulViMotion and DeepMesh. Details about the modified baselines are provided in Appendix C.
@@
Figure 3: 1-Slice Comparison. (Left) We predict meshes by deforming from ED to ES frame. Color indicates the magnitude of point-to-surface error. (Right) Segmentation results on slices at the apex and base location. Note 4DMR can only output motion for myocardium.
@@
Figure 4: Motion sequence prediction results using 1-/Full-slices by deforming from ED frame to the target frame. Corresponding Volume-frame curve is given on the right. #1: NOR, #2: DCM.
@@ 4.2 Motion Estimation from a Few Slices
We present results under the online few-slice scenario that represents the primary motivation of this work. Fig. 3 shows a representative motion reconstruction using only the central slice as input: one from a normal (NOR) subject and another from a case of dilated cardiomyopathy (DCM). Due to stronger contractile function, the normal heart exhibits larger deformations, making motion prediction inherently more challenging, as reflected by higher reconstruction errors. Nevertheless, our method delivers accurate predictions in both cases. Moreover, segmentation results on slices near the apex and base demonstrate that our model produces reliable deformations even in regions distant from the input slice. Complete reconstructed sequences and their corresponding volume–time curves are shown in Fig. 4.
Quantitative results on the full test set are reported in Tab. 1 and the top section of Tab. 2, including Chamfer Distance, Dice scores, and cardiac function indices. For consistency, all experiments were performed using the central 1 or 5 slices, though slices at arbitrary positions could have been used, as discussed in Section 4.4. TetHeart consistently outperforms all baselines without requiring ground-truth contours or segmentations as input. It infers cardiac motion directly from raw CMR images, eliminating the need for time-consuming test-time optimization, and achieves an inference speed of 12 FPS on an NVIDIA V100 GPU. Further acceleration could be achieved through model quantization or lightweight encoders to meet strict real-time constraints. Notably, even when provided with a single slice, TetHeart produces meaningful physiological indices, underscoring its suitability for clinical deployment.
Although Ours-Mesh and Ours-SDF incorporate the same sparse feature handling strategy as our full model, TetHeart still achieves superior performance. This confirms that the tetrahedral representation effectively preserves spatial coherence and serves as a more expressive basis for 3D motion inference under sparse observations. Fig. 5 illustrates performance as a function of slice count: the fewer slices are used, the larger the gap between TetHeart and the ablated variants. Nevertheless, both variants still outperform all prior methods by a clear margin, demonstrating that the AFA module and the training strategy form a robust and generalizable solution for incomplete observations.
The relatively poor performance of 4DMR can be attributed to its simple MLP architecture, which may lack the representational power to generalize well to the large and heterogeneous datasets considered here. Moreover, its test-time optimization strategy does not generalize well to severely under-sampled inputs. The weaker performance of MR-Net likely arises from the lack of reconstruction features and its sole reliance on contour points for deformation prediction, which overlooks rich cine image information.
4.3 Motion Estimation from Full Stacks
@@
Since we had to modify existing techniques to compare them to ours in the few-slice scenario, we now compare our method to the original baselines on the full set of slices they were designed to operate on and report the results in Tab. 3 and the bottom of Tab. 2. In Fig. 4 and S2, we show qualitative results on the same hearts as in Fig. 3, but now using all slices as input instead of only one.
Remarkably, although the main driver of this work was operating on as few slices as possible, we also consistently outperform the baselines in the full-slice setting. We attribute this to several factors. Using predicted contours in MR-Net leads to slightly worse performance, indicating such a two-stage automated pipeline is prone to error accumulation, consistent with the conclusions in the original paper [10]. MulViMotion estimates voxel-wise motion fields and reconstructs meshes from warped segmentation maps, which is inherently less accurate than directly modeling meshes. Compared to DeepMesh, our model benefits from a shared encoding network, allowing the deformation branch to leverage features from the reconstruction branch. Moreover, the AFA module dynamically aggregates spatially adjacent features, improving prediction quality.
4.4 Generalization without Retraining
We now perform external evaluation to test the generalizability of our model to the iCMR and 4DM datasets without retraining. In Fig. 6, we show sequence predictions on iCMR dataset with the spatial position of the real-time captured 2D slice visualized. As can be seen, TetHeart reconstructs plausible heart motion across the cardiac cycle under both rest and exercise situation. Moreover, the predicted volume–time curves confirm the physiological plausibility of our motion reconstruction. In the exercise condition, the heart exhibits a shorter ES–ED recovery period and a reduced duration of each cardiac cycle, aligns well with expectation. Please refer to Appendix E for more details and the external evaluation results on the 4DM dataset.
Figure 6: Motion sequence prediction on the iCMR dataset. (Left) Spatial configuration of the initial SAX stack and the online observations. (Middle) Cardiac motion sequences for each subject under rest and exercise conditions. (Right) Volume–time curves.
4.5 Ablation Studies
We conduct ablation studies to evaluate our design choices. First, we examine the AFA module. Simply stacking 2D features yields similar performance to AFA in the full-slice setting but fails to generalize when the number or positions of slices vary, highlighting the need for a flexible aggregation mechanism. Removing position encodings moderately degrades performance, confirming the importance of spatial information, while using only position encodings leads to a substantial drop—showing that both image features and spatial coordinates are essential for effective motion prediction. When the motion branch is trained without parameter sharing and randomly initialized, performance decreases significantly, demonstrating that our shared-parameter design effectively transfers 3D knowledge from reconstruction to motion modeling.
Lastly, we ablate the impact of the distillation loss. Without it, the model performs slightly better in the full-slice setting but encodes less discriminative features from sparse 2D inputs, resulting in poorer few-slice performance. Combining the distillation loss with our aggressive sampling strategy consistently improves results across all scenarios, underscoring the importance of this training scheme for improving robustness and generalization. Additional ablation results are provided in Appendix F.
  Method M&Ms Myocardium CD (mm2) $\downarrow$
1-Slice 5-Slice Full-Slice
Ours 9.76 7.93 4.11
- AFA module - - 4.22
- position embedding 11.23 9.04 4.67
- image feature in query $Q$ 16.67 12.37 6.44
- shared encoding network 20.66 11.63 6.43
- distillation loss 10.34 8.12 4.03
Table 4: Ablation study results on M&Ms dataset.
5 Conclusion
We have presented TetHeart, a unified framework for 4D cardiac shape and motion reconstruction from 2D CMR slices, operating robustly from full stacks down to a single real-time slice. For the first time, we show that accurate 3D cardiac motion can be inferred from sparse intra-operative CMR data, enabling real-time tracking and guidance during interventions. This has the potential to shift cardiac reconstruction from retrospective analysis to live clinical application, laying the foundation for patient-specific digital twins of the beating heart and paving the way for adaptive, predictive, and personalized image-guided procedures.
As an immediate next step, we plan to deploy TetHeart in an interventional MRI suite at our collaborating hospital. Looking ahead, we aim to extend this framework to other dynamic organs, integrate multimodal imaging, and develop predictive digital twins that anticipate motion under both physiological and interventional conditions—advancing toward a new generation of real-time, personalized, image-guided medicine.
Supplementary Material
Appendix A Datasets and Metrics
A.1 Unified Dataset
ACDC features CMR sequences of about 30 frames for 100 patients, each one covering at least one cardiac cycle and with a slice thickness of 5-10 mm. M&Ms and M&Ms-2 are made of multi-center and multi-vendor collected CMR sequences for 375 and 360 subjects, respectively. While featuring similar resolutions and sequence lengths as ACDC dataset, these two datasets cover a broader range of cardiac pathologies. For these datasets, segmentation masks for the left ventricle (LV), right ventricle (RV), and myocardium (MYO) are provided, but only for the end-diastolic (ED) and end-systolic (ES) frames. As in earlier studies [28, 52, 48, 33], we register and fit a Statistic Shape Model (SSM) [15, 2] to segmentation mask at the ED and ES frames using non-rigid image registration methods to generate cardiac meshes.
A.2 iCMR Dataset
The iCMR dataset was collected at the Center for Interventional MRI of our collaborating university hospital. For each subject, a stack of short-axis (SAX) breath-hold cine slices was acquired at rest on a 1.5T interventional system (Siemens Healthineers, Magnetom Aera) using an accelerated steady-state free precession (SSFP) sequence covering the entire left ventricle. In addition, single SAX cine slices were acquired at mid-ventricular levels with slight angulations relative to the full SAX stack. Because both heart shape and rate may vary during intervention, updating of the 3D heart model is often required to ensure precise intracardiac manipulations. To mimic heart-rate changes, five volunteers exercised inside the interventional MRI (bending and extending their legs in two parallel plastic gutters for 2–3 minutes), thereby increasing their heart rates by approximately 20–30 beats per minute. Once elevated, single SAX cine slices were reacquired at the same mid-ventricular levels as during rest. For real-time acquisition, the online sequences had lower spatial resolution than the offline short-axis CMR stacks to meet the latency constraint.
A.3 Evaluation Metrics
For evaluation, we compare the predicted 3D mesh obtained by deforming the ED frame to the ES frame with the ground-truth 3D mesh at the ES frame. We use L2-norm Chamfer Distance (CD) to measure 3D reconstruction quality, which is calculated as:
$\mathrm{CD}(P,Q)=\frac{1}{|P|}\sum_{p\in P}\min_{q\in Q}\|p-q\|_{2}^{2}+\frac{1}{|Q|}\sum_{q\in Q}\min_{p\in P}\|q-p\|_{2}^{2}.$
Dice score is also reported to quantify the intra-slice segmentation results, calculated as:
$\mathrm{DSC}(A,B)=\frac{2|A\cap B|}{|A|+|B|}.$
Besides these quality measures, we also report the Mean-Absolute Error (MAE) of LV/RV ES Volume (LVESV/RVESV), LV/RV ejection fraction (LVEF/RVEF) derived from our predicted meshes. The ejection fraction is calculated as
$XEF=\frac{XEDV-XESV}{XEDV},X\in\{LV,RV\},$
here EDV denotes the volume of the ED meshes. For our method, we use the predicted ED meshes from the reconstruction branch for calculation. For baselines require ground-truth input meshes, we directly use the input meshes for calculation.
The MAE of ES volume and ejection fraction is then calculated as
$MAE_{esv}=|ESV_{p}-ESV_{gt}|,~MAE_{ef}=|EF_{p}-EF_{gt}|.$
The ground-truth volume and ejection fraction are derived from the ground-truth meshes. LVEF and RVEF are representative cardiac function indexes which could be used for clinical diagnosis.
Appendix B Implementation Details
Figure S1: Motion sequence prediction results using 1-/Full-slices by deforming from ED frame to the target frame. Corresponding Volume-frame curve is given on the right. #1: NOR, #2: DCM.
Appendix C Baselines
We compare our method against several existing approaches for 4D shape reconstruction. For the online interventional setting, we adapted several baselines so that they can operate under sparse inputs. 4DMR [52] is an implicit representation method that reconstructs myocardium mesh sequence by optimizing decoupled shape and motion latent codes at test time, minimizing the discrepancy between the predicted mesh and ground-truth contour points. We adapt it by restricting the optimization to incomplete contours extracted from the available few slices, while still using full contours for the initial frame. MR-Net [10] is a template-mesh-based approach that takes ground-truth contour points as input and leverages a hybrid PointNet + 3D CNN architecture to predict vertex deformations. Since MR-Net directly encodes point sets, it can naturally accommodate incomplete inputs, a scenario already considered in the original paper. Although originally designed for static reconstruction, we adapt it for motion estimation by setting the ED-frame mesh as the template.
Figure S2: Full-Slice Comparison. (Left) We predict 3D meshes by deforming from ED to ES frame. Color indicates the magnitude of point-to-surface error in $m$. (Right) Segmentation results on slices at the apex and base location. Note 4DMR and DeepMesh can only output for myocardium.